# Dynamic Asset Allocation with Asset-Specific Regime Forecasts

### v1

@@ Abstract
This article introduces a novel hybrid regime identification-forecasting framework designed to enhance multi-asset portfolio construction by integrating asset-specific regime forecasts. Unlike traditional approaches that focus on broad economic regimes affecting the entire asset universe, our framework leverages both unsunpervised and supervised learning to generate tailored regime forecasts for individual assets. Initially, we use the statistical jump model, a robust unsupervised regime identification model, to derive regime labels for historical periods, classifying them into bullish or bearish states based on features extracted from an asset return series. Following this, a supervised gradient-boosted decision tree classifier is trained to predict these regimes using a combination of asset-specific return features and cross-asset macro-features. We apply this framework individually to each asset in our universe. Subsequently, return and risk forecasts which incorporate these regime predictions are input into Markowitz mean-variance optimization to determine optimal asset allocation weights. We demonstrate the efficacy of our approach through an empirical study on a multi-asset portfolio comprising twelve risky assets, including global equity, bond, real estate, and commodity indexes spanning from 1991 to 2023. The results consistently show outperformance across various portfolio models, including minimum-variance, mean-variance, and naive-diversified portfolios, highlighting the advantages of integrating asset-specific regime forecasts into dynamic asset allocation.
@@ 1 Introduction
The cyclical regime-switching behavior of asset returns is well-documented across various asset classes, including equities (Hardy, , 2001), fixed income (Gray, , 1996), commodities (Choi and Hammoudeh, , 2010), currencies (Reus and Mulvey, , 2016), and cryptocurrencies (Ardia et al., , 2019). This behavior manifests in patterns such as bullish or bearish markets (Pagan and Sossounov, , 2003), periods of high or low volatility (Schwert, , 1989), and risk on or off episodes (De Bock and de Carvalho Filho, , 2015). A specific regime denotes extended, consecutive periods that display homogeneous market dynamics regarding returns, volatility, and sentiment, among other factors, while a regime shift signifies an abrupt yet persistent change in market behavior. Regime-switching models are highly valued for their interpretability, as the regimes they identify often correlate with real-world events like different phases of the business cycle (Hamilton, , 1989) or shifts in monetary policy (Ang and Bekaert, , 2002). These models not only align with changing fundamentals that can often be interpreted only ex-post, but also facilitate ex-ante real-time forecasting (Ang and Timmermann, , 2012). Notably, Ang and Bekaert, (2004) was among the first to empirically demonstrate the benefits of incorporating regimes into asset allocation in an out-of-sample setting.
Previous studies on regime-based asset allocation typically focus on identifying or predicting a limited number (usually 2–4) of economic regimes, such as expansion and recession periods, and subsequently employ regime-dependent strategies. These economic regimes are often inferred from macroeconomic indicators (Kim and Kwon, , 2023; Mueller-Glissmann and Ferrario, , 2024), or derived from a major equity index using Markov-switching models (Nystrup et al., , 2015). Regime-dependent actions might involve tilting weights towards favorable assets under each regime (Nystrup et al., , 2018), employing dynamic risk budgets (James et al., , 2019), or integrating regime-dependent return and covariance forecasts into portfolio optimization (Uysal and Mulvey, , 2021). This approach presumes that a few broad economic regimes can explain the variation in time-varying asset behaviors uniformly across the entire asset universe – for instance, stocks typically underperform during recessions, while bonds may perform oppositely. While the methodology’s strength lies in its ability to make asset performance interpretable in terms of the macro-environment, it may be overly optimistic to believe that such a simplified set of regime concepts can adequately capture the complex behaviors of a diverse range of assets.
@@ 3.2 Regime Identification
In our regime identification approach, we prioritize two key requirements: first, the identified regimes must be able to directly enhance the portfolio performance of asset allocation; second, these regimes should enable accurate forecasting in the subsequent step. These priorities guide our shift to identifying asset-specific market regimes using jump models, offering several benefits. Firstly, we focus on market dynamics over broad economic conditions to more closely link identified regimes with the ultimate portfolio performance. To emphasize this focus, we refer to our identified regimes as market regimes rather than economic regimes. Secondly, by identifying regimes for each asset individually within the universe, we capture the nuances of every market, avoiding the simplification of using a major equity market like US LargeCap to represent all asset classes. This targeted approach also mitigates the risk of incorrect regime forecasts by employing an ensemble of them. Indeed, the consequences of incorrect regime forecasts can be severe; as Hess, (2006) notes, “a wrong regime forecast may not only lead to a non-optimal but to a detrimental allocation in the contrary direction”. The third benefit arises from the fact that our regime forecasts, being asset-specific, are somewhat analogous to directly forecasting all asset returns. This similarity leads to a significant advantage: our identified regimes typically exhibit an enhanced signal-to-noise ratio (SNR) compared to direct return forecasting, thereby creating favorable conditions for the subsequent forecasting step. This improvement stems from the persistence inherent in the regime definition, leading to regimes that effectively act as a heavily smoothed version of raw returns, for the purpose of forecasting121212 Preprocessing raw returns in defining forecasting targets proves efficient among practitioners for enhancing SNR. For instance, removing return components explained by factors, as noted by Kaniel et al., (2023), leads to “higher accuracy and better portfolio performance” when predicting abnormal returns instead of total returns for mutual funds. Similarly, smoothing raw returns over a forward window, as implemented in Aït-Sahalia et al., (2022), results in a “less noisy response variable”, which is “more indicative of aggregated trading behavior [in a forward window] rather than at an arbitrary point in the near future.” . Predictability varies across assets, and we fine-tune the jump penalty, which moderates the SNR, as outlined in Section 3.4, to optimize performance for each asset.
@@ 3.2.1 Statistical Jump Models
Our features used in JMs include eight financial measures calculated from an individual asset excess return series, designed to determine whether the asset is in a bullish or bearish market. Referred to as return features, these measures are asset-specific since a JM is fitted individually. Specifically we utilize exponentially weighted moving (EWM) downside deviation (DD) in log scale, average returns, and Sortino ratio across various halflives, as summarized in Table 2. We prefer DD to volatility as a risk measure due to its emphasis on downside risk, and we apply a log transformation to DD measures to ensure stable model fitting, especially as volatility measures tend to spike during market turmoil. To alleviate feature correlation, we select only two halflives for DD since risk measures display much higher correlation across different halflives compared to return measures161616 We exclude the two DD features for three assets: AggBond, Treasury, and Gold, as preliminary in-sample analysis indicates that DD features do not distinctly separate under the two fitted regimes. However, the forecasting algorithm described in Section 3.3 includes all eight return features for all assets, thanks to its improved accommodation to high-dimensionality. . The EWM Sortino ratio, defined as the ratio of EWM average returns over EWM DD, represents risk-adjusted returns. For exponential smoothing, we choose halflives corresponding to one week, two weeks, and a month in trading days, avoiding shorter halflives due to increased noise.
@@ 3.4 Hyperparameter Tuning
First and foremost, we prioritize financial metrics for our selection criterion over statistical metrics such as classification accuracy used by Uysal and Mulvey, (2021) or information criteria used by Cortese et al., (2024). Since regime labels are identified by our method, they may not accurately represent the “ground truth” necessary for calculating classification accuracy. On the contrary, while the 0/1 strategy may appear too simplistic for real-world portfolio management, it effectively evaluates the financial implication of the statistical accuracy of the regime forecasts for a specific asset. We use the Sharpe ratio as our primary selection criterion, noting that typically the optimal $\lambda$ excels across various (risk-adjusted) return metrics, including cumulative return, Sharpe ratio, Sortino ratio, and information ratio. Throughout this article, the performance of the 0/1 strategy serves to evaluate financially the prediction accuracy of our regime forecasts, as presented in Section 5.1.
@@ 4.2 Minimum-Variance Portfolio
MinVar portfolio, identified as the leftmost point on the efficient frontier given risk on the $x$-axis, offers the practical advantage of not requiring return forecasts, which are notoriously difficult to estimate accurately. Instead, it relies on the more reasonably estimated covariance matrix, often refined through factor models and shrinkage methods. More recently, MinVar strategy has been recognized as a market anomaly. Ang et al., (2006) observed that stocks with higher historical idiosyncratic volatility tend to yield lower realized returns, and Clarke et al., (2006) noted that MinVar portfolio effectively reduces risk without compromising returns, leading to an improved risk-adjusted return compared to the market. Nowadays, the growth of exchange-traded funds (ETFs) has democratized access to sophisticated optimization-based investment strategies for retail investors. For instance, the ETF USMV, which applies a minimum volatility strategy to over a hundred US equities, had a market value exceeding 24 billion dollars at the time of writing.
@@ 4.4 Equally-Weighted Portfolio
EW portfolio, which distributes capital equally across a universe of assets, serves as a notable example of a naive diversification rule that does not depend on any forecasting or optimization skill. Despite debates sparked by a series of articles including DeMiguel et al., (2009); Kirby and Ostdiek, (2012); Behr et al., (2013), there is a consensus that, if not properly implemented, MVO-based portfolio methods including MinVar and MV, can easily underperform simple diversification approaches like EW in an out-of-sample setting, despite their theoretical soundness. Beyond its frequent use as a benchmark in academic research, the practical implementation of EW strategies has a rich history, ranging from the first index fund established in 1971 based on an equally-weighted index of New York Stock Exchange equities, to the popularity of ETFs such as RSP, which applies equal weights to the stocks in the S&P 500 Index191919 Based on a speech by John C. Bogle about the origin of indexing, retrieved from https://boglecenter.net/wp-content/uploads/Superbowl-12-5-01.pdf. The 1971 fund was designed for a pension, while the fund Bogle created is recognized as the first index mutual fund.
RSP had a market value of over 55 billion dollars at the time of writing.
@@ 5 Empirical Results
This section presents the empirical results of strategy performance. In Section 5.1, we detail the performance of the 0/1 regime-switching strategy applied to each individual asset, aimed at assessing the financial implication of the accuracy of our asset-specific regime forecasts. Section 5.2 then presents the performance of dynamic asset allocation strategies that incorporate our regime forecasts through three distinct portfolio models: minimum-variance (MinVar), mean-variance (MV), and equally-weighted (EW) portfolios. A sensitivity analysis follows to evaluate the robustness of these results. Throughout the empirical study, a one-way transaction cost of 5 basis points is applied.
@@ 5.1 Regime-Switching Strategy Performance for Individual Assets
In summary, the financial significance of the regime forecasts generated by both the JM and JM-XGB approach is clear, highlighting the potential benefits of integrating these forecasts into practical portfolio management. While the comparison to the JM approach might not be fair enough – given the lack of fine-tuning and the exclusion of macro-features in the JM approach – the advantages of using a combined regime identification-forecasting framework becomes evident. Therefore, the next subsection will focus exclusively on displaying the portfolio performance based on the JM-XGB approach.

### v2

@@ Abstract
This article introduces a novel hybrid regime identification-forecasting framework designed to enhance multi-asset portfolio construction by integrating asset-specific regime forecasts. Unlike traditional approaches that focus on broad economic regimes affecting the entire asset universe, our framework leverages both unsupervised and supervised learning to generate tailored regime forecasts for individual assets. Initially, we use the statistical jump model, a robust unsupervised regime identification model, to derive regime labels for historical periods, classifying them into bullish or bearish states based on features extracted from an asset return series. Following this, a supervised gradient-boosted decision tree classifier is trained to predict these regimes using a combination of asset-specific return features and cross-asset macro-features. We apply this framework individually to each asset in our universe. Subsequently, return and risk forecasts which incorporate these regime predictions are input into Markowitz mean-variance optimization to determine optimal asset allocation weights. We demonstrate the efficacy of our approach through an empirical study on a multi-asset portfolio comprising twelve risky assets, including global equity, bond, real estate, and commodity indexes spanning from 1991 to 2023. The results consistently show outperformance across various portfolio models, including minimum-variance, mean-variance, and naive-diversified portfolios, highlighting the advantages of integrating asset-specific regime forecasts into dynamic asset allocation.
@@ 1 Introduction
The cyclical regime-switching behavior of asset returns is well-documented across various asset classes, including equities (Hardy, , 2001), fixed income (Gray, , 1996), commodities (Choi and Hammoudeh, , 2010), currencies (Reus and Mulvey, , 2016), and cryptocurrencies (Ardia et al., , 2019). This behavior manifests itself in patterns such as bullish or bearish markets (Pagan and Sossounov, , 2003), periods of high or low volatility (Schwert, , 1989), and risk on or off episodes (De Bock and de Carvalho Filho, , 2015). A specific regime denotes extended, consecutive periods that display homogeneous market dynamics regarding returns, volatility, and sentiment, among other factors, while a regime shift signifies an abrupt yet persistent change in market behavior. Regime-switching models are highly valued for their interpretability, as the regimes they identify often correlate with real-world events like different phases of the business cycle (Hamilton, , 1989) or shifts in monetary policy (Ang and Bekaert, , 2002). These models not only align with changing fundamentals that can often be interpreted only ex-post, but also facilitate ex-ante real-time forecasting (Ang and Timmermann, , 2012). Notably, Ang and Bekaert, (2004) was among the first to empirically demonstrate the benefits of incorporating regimes into asset allocation in an out-of-sample setting.
Previous studies on regime-based asset allocation typically focus on identifying or predicting a limited number (usually 2–4) of economic regimes, such as expansion and recession periods, and subsequently employ regime-dependent strategies. These economic regimes are often inferred from macroeconomic indicators (Kim and Kwon, , 2023; Mueller-Glissmann and Ferrario, , 2024), or derived from a major equity index using Markov-switching models (Nystrup et al., , 2015). Regime-dependent actions might involve tilting weights towards favorable assets under each regime (Nystrup et al., , 2018), employing dynamic risk budgets (James et al., , 2019), or integrating regime-dependent return and covariance forecasts into portfolio optimization (Uysal and Mulvey, , 2021). This approach presumes that a few broad economic regimes can explain the variation in time-varying asset behaviors uniformly across the entire asset universe – for instance, stocks typically underperform during recessions, while bonds may outperform. While the methodology’s strength lies in its ability to make asset performance interpretable in terms of the macro-environment, it may be overly optimistic to believe that such a simplified set of regime concepts can adequately capture the complex behaviors of a diverse range of assets.
@@ 3.2 Regime Identification
In our regime identification approach, we prioritize two key requirements: first, the identified regimes must be able to directly enhance the portfolio performance of asset allocation; second, these regimes should enable accurate forecasting in the subsequent step. These priorities guide our shift to identifying asset-specific market regimes using jump models, offering several benefits. Firstly, we focus on market dynamics over broad economic conditions to more closely link identified regimes with the ultimate portfolio performance. To emphasize this focus, we refer to our identified regimes as market regimes rather than economic regimes. Secondly, by identifying regimes for each asset individually within the universe, we capture the nuances of every market, avoiding the simplification of using a major equity market like US LargeCap to represent all asset classes. This targeted approach also mitigates the risk of incorrect regime forecasts by employing an ensemble of them. Indeed, the consequences of incorrect regime forecasts can be severe; as Hess, (2006) notes, “a wrong regime forecast may not only lead to a non-optimal but to a detrimental allocation in the contrary direction”. The third benefit arises from the fact that our regime forecasts, being asset-specific, are somewhat analogous to directly forecasting all asset returns. This similarity leads to a significant advantage: our identified regimes typically exhibit an enhanced signal-to-noise ratio (SNR) compared to direct return forecasting, thereby creating favorable conditions for the subsequent forecasting step. This improvement stems from the persistence inherent in the regime definition, leading to regimes that effectively act as a heavily smoothed version of raw returns, for the purpose of forecasting121212 Preprocessing raw returns in defining forecasting targets proves efficient for enhancing SNR. For instance, removing return components explained by factors, as noted by Kaniel et al., (2023), leads to “higher accuracy and better portfolio performance” when predicting abnormal returns instead of total returns for mutual funds. Similarly, smoothing raw returns over a forward window, as implemented in Aït-Sahalia et al., (2022), results in a “less noisy response variable”, which is “more indicative of aggregated trading behavior [in a forward window] rather than at an arbitrary point in the near future.” . Predictability varies across assets, and we fine-tune the jump penalty, which moderates the SNR, as outlined in Section 3.4, to optimize performance for each asset.
@@ 3.2.1 Statistical Jump Models
Our features used in JMs include eight financial measures calculated from an individual asset excess return series, designed to determine whether the asset is in a bullish or bearish market. Referred to as return features, these measures are asset-specific since a JM is fitted individually. Specifically we utilize exponentially weighted moving (EWM) downside deviation (DD) in log scale, average returns, and Sortino ratio across various halflives, as summarized in Table 2. We prefer DD to volatility as a risk measure due to its emphasis on downside risk, and we apply a log transformation to DD measures to ensure stable model fitting, especially as volatility measures tend to spike during market turmoil. To alleviate feature correlation, we select only two halflives for DD since risk measures display much higher correlation across different halflives compared to return measures161616 We exclude the two DD features for three assets: AggBond, Treasury, and Gold, as preliminary in-sample analysis indicates that risk features do not distinctly separate under the two fitted regimes. However, the forecasting algorithm described in Section 3.3 includes all eight return features for all assets, thanks to its improved accommodation to high-dimensionality. . The EWM Sortino ratio, defined as the ratio of EWM average returns over EWM DD, represents risk-adjusted returns. For exponential smoothing, we choose halflives corresponding to one week, two weeks, and a month in trading days, avoiding shorter halflives due to increased noise.
@@ 3.4 Hyperparameter Tuning
First and foremost, we prioritize financial metrics for our selection criterion over statistical metrics such as classification accuracy used by Uysal and Mulvey, (2021) or information criteria used by Cortese et al., (2024). Since regime labels are identified by our method, they may not accurately represent the “ground truth” necessary for calculating classification accuracy. On the contrary, the 0/1 strategy effectively evaluates the financial implication of the statistical accuracy of the regime forecasts for a specific asset. A comprehensive discussion of the 0/1 strategy and its effectiveness in mitigating downside risk in asset management is detailed in Shu et al., (2024). We use the Sharpe ratio of this strategy as our primary selection criterion, noting that typically the optimal $\lambda$ excels across various (risk-adjusted) return metrics, including cumulative return, Sharpe ratio, Sortino ratio, and information ratio. Throughout this article, the performance of the 0/1 strategy serves to evaluate financially the prediction accuracy of our regime forecasts, as presented in Section 5.1.
@@ 4.2 Minimum-Variance Portfolio
MinVar portfolio, identified as the leftmost point on the efficient frontier given risk on the $x$-axis, offers the practical advantage of not requiring return forecasts, which are notoriously difficult to estimate accurately. Instead, it relies on the more reasonably estimated covariance matrix, often refined through factor models and shrinkage methods. More recently, MinVar strategy has been recognized as a market anomaly. Ang et al., (2006) observed that stocks with higher historical idiosyncratic volatility tend to yield lower realized returns, and Clarke et al., (2006) noted that MinVar portfolio effectively reduces risk without compromising returns, leading to an improved risk-adjusted return compared to the market. Nowadays, the growth of exchange-traded funds (ETFs) has democratized access to sophisticated optimization-based investment strategies for retail investors. For instance, the ETF USMV, which employs a minimum volatility strategy across a broad range of US equities, had a market value exceeding 24 billion dollars at the time of writing.
@@ 4.4 Equally-Weighted Portfolio
EW portfolio, which distributes capital equally across a universe of assets, serves as a notable example of a naive diversification rule that does not depend on any forecasting or optimization skill. Despite debates sparked by a series of articles including DeMiguel et al., (2009); Kirby and Ostdiek, (2012); Behr et al., (2013), there is a consensus that, if not properly implemented, MVO-based portfolio methods including MinVar and MV, can easily underperform simple diversification approaches like EW in an out-of-sample setting, despite their theoretical soundness. Beyond its frequent use as a benchmark in academic research, the practical implementation of EW strategies has a rich history, ranging from the first index fund established in 1971 based on an equally-weighted index of New York Stock Exchange equities, to the popularity of ETFs such as RSP, which applies equal weights to the stocks in the S&P 500 Index191919 Based on a speech by John C. Bogle about the origin of indexing, retrieved from https://boglecenter.net/wp-content/uploads/Superbowl-12-5-01.pdf. The 1971 fund was designed for a pension, while the fund Bogle created is recognized as the first index mutual fund. RSP had a market value of over 55 billion dollars at the time of writing. .
@@ 5 Empirical Results
This section presents the empirical results of strategy performance over the testing period from 2007 to 2023. In Section 5.1, we detail the performance of the 0/1 regime-switching strategy applied to each individual asset, aimed at assessing the financial implication of the accuracy of our asset-specific regime forecasts. Section 5.2 then presents the performance of dynamic asset allocation strategies that incorporate our regime forecasts through three distinct portfolio models: minimum-variance (MinVar), mean-variance (MV), and equally-weighted (EW) portfolios. A sensitivity analysis follows to evaluate the robustness of these results. Throughout the empirical study, a one-way transaction cost of 5 basis points is applied.
@@ 5.1 Regime-Switching Strategy Performance for Individual Assets
In summary, the financial significance of the regime forecasts generated by both the JM and JM-XGB approach is clear, highlighting the potential benefits of integrating these forecasts into practical portfolio management. These advantages are consistent with previous empirical findings, such as those reported in Shu et al., (2024); Bosancic et al., (2024), which demonstrate the value of leveraging a regime-switching signal via the 0/1 strategy. While the comparison to the JM approach might not be fair enough – given the lack of fine-tuning and the exclusion of macro-features in the JM approach – the advantages of using a combined regime identification-forecasting framework becomes evident. Therefore, the next subsection will focus exclusively on displaying the portfolio performance based on the JM-XGB approach.